# A Centralized Planning and Distributed Execution Method for Shape Filling with Homogeneous Mobile Robots

### v1

@@ Abstract
Abstract: Nature has inspired humans in different ways. The formation behavior of animals can perform tasks that exceed individual capability. For example, army ants could transverse gaps by forming bridges, and fishes could group up to protect themselves from predators. The pattern formation task is essential in a multiagent robotic system because it usually serves as the initial configuration of downstream tasks, such as collective manipulation and adaptation to various environments. The formation of complex shapes, especially hollow shapes, remains an open question. Traditional approaches either require global coordinates for each robot or are prone to failure when attempting to close the hole due to accumulated localization errors. Inspired by the ribbon idea introduced in the additive self-assembly algorithm by the Kilobot team, we develop a two-stage algorithm that does not require global coordinates information and effectively forms shapes with holes. In this paper, we investigate the partitioning of the shape using ribbons in a hexagonal lattice setting and propose the add-subtract algorithm based on the movement sequence induced by the ribbon structure. This advancement opens the door to tasks requiring complex pattern formations, such as the assembly of nanobots for medical applications involving intricate structures and the deployment of robots along the boundaries of areas of interest. We also provide simulation results on complex shapes, an analysis of the robustness as well as a proof of correctness of the proposed algorithm.
@@ 1 Introduction
Multi-Robot Systems (MRS), as suggested by the name, involves the study of groups of heterogeneous or homogeneous autonomous robots tackling a wide range of tasks with varying characteristics in the same environment [1]. Among the various tasks assigned to a group of autonomous robots, the deployment task has received significant attention since it is usually used to set up the initial configuration for other downstream tasks such as exploration, surveillance and cooperative manipulation. The pattern formation task, as a variant of the robot deployment tasks, aims to assemble a collection of robots into proper positions that resemble a specific pattern when viewed globally [2]. Multiple researches have been performed on the pattern formation using homogeneous autonomous robots.
In [3], Stoy and Kasper proposed a shape formation algorithm utilizing a porous scaffold and established two types of gradients: the wander gradient, which is generated from wandering robots outside the desired shape, and the hole gradient, which is generated from the unfilled spaces inside the desired shape. A wandering robot descends the hole gradient and ascends the wander gradient, resulting in the robots inside the shape moving away from the boundary and propagating the unfilled spaces towards the boundary. In such a way, this algorithm is able to reconfigure a random shape into a desired non-hollow shape.
@@
In [10], Derakhshandeh et al. designed a shape formation algorithm based on the triangular expansion behavior of programmable matters. The algorithm can be adapted to any complex shape composed of equilateral triangles. However, similar to Yang’s approach, without global information on coordinates, the robots may not be able to form the holes correctly due to accumulated localization errors.
In this article, we propose a two-stage pattern formation algorithm, the add-subtract algorithm, to handle formation shapes with holes. The algorithm is inspired by the ribbon idea introduced in the additive self-assembly algorithm [4]. In our algorithm, we assume a grid-based workspace and choose the hexagonal lattice due to its compactness and suitability for disk-shaped robots. Several studies have explored hexagonal lattice configurations. K. Tomita et al. developed a nucleation method using hexagon-shaped robots [11]; S. Prabhu et al. investigated hexagonal formations using artificial force fields [12]; and M. Rubenstein et al. utilized hexagonal arrangements to organize idle robots [4]. Building on this foundation, we propose a systematic method to partition the desired shape into layers (referred to as ribbons) and plan the path for each individual robot.
Our method is different from others in terms of the following aspects:
@@
    unlike the original additive self-assembly algorithm, whose outcome is not known in advance due to the lack of planning. Our algorithm leverages the geometry of the hexagonal lattice to specify the boundary smoothness of the desired shape and the correctness of our algorithm can be guaranteed; and
@@
    unlike most other existing algorithms that can only form a simple connected shape, by adding the subtractive stage, our algorithm is able to form shapes with multiple holes; and
@@
    unlike existing pattern formation algorithms designed for complex hollow shapes, our algorithm does not rely on the global coordinates of each robot, thus, facilitate effective distributed execution.
The remainder of the paper is organized as follows. First, the problem is introduced and formulated in section 2. The add-subtract algorithm is introduced and analyzed in section 3. To ensure a smooth presentation flow, while we present the outline of the proof of correctness in section 4, details of all proofs are presented in the Appendices. Simulation results are given in section 5. Failure handling is discussed in section 6, and conclusions are drawn in section 7 together with a discussion of future research directions. For the pseudocode of the algorithm, please refer to appendix A. For the proof of the properties of the ribbon structure and the proof of correctness of the algorithm can be found in appendix B and appendix C.
@@ 2.1 Problem Description
In this paper we study the problem of filling up a given shape with a given collection of small mobile robots. The shape may contain holes, and each mobile robot has a limited localization capability, forcing them to rely on neighboring robots to precisely determine its current location. Due to this limitation, during the shape-filling process, robots must follow a certain releasing sequence to ensure that each newly released robot can rely on previously released and positioned robots to precisely locate itself, before moving into a designated position in the shape. We assume that each robot has sufficient memory to store several pieces of critical information such as the releasing sequence, which encoded boundaries of the shape and internal holes. In addition, we assume that each robot can only communicate with robots within a neighborhood whose maximum radius is known in advance. Next, we provide more details of the system setup.
@@
We consider a user-specified shape $S_{0}$ defined in Euclidean coordinate system in $\mathbb{R}^{2}$ which is topologically homeomorphic to an open disk in $\mathbb{R}^{2}$. We consider a set $D_{0}$ of user-specified Holes defined in Euclidean coordinate system in $\mathbb{R}^{2}$, each of which is topologically homeomorphic to a collection of closed disks in $\mathbb{R}^{2}$ and contained inside $S_{0}$. With a slight abuse of notation, we use $D_{0}$ to denote the union of all such holes. Thus, $D_{0}$ is a subset of $S_{0}$.
@@
    all the lattice points in $S$ are connected (in the sense that there is a path between any pair of lattice point in $S$ through adjacent lattice points in $S$); and
@@
    For each prefix subsequence $(a_{1},b_{1},1),\ldots,(a_{k},b_{k},k)$, $1\leq k\leq M$, let $A_{k}=\{a_{1},\ldots,a_{k-1}\}$, $B_{k}=\{b_{1},\ldots,b_{k-1}\}$, and $\overline{B_{k}}\subseteq B_{k}$ be the set of stopping positions that have been reached by the corresponding moving robots. Then the starting position of the $k$th robot $a_{k}\in I\oplus\overline{B_{k}}\circleddash A_{k}$ and the robot can only relay on robots centered at $I\oplus\overline{B_{k}}\circleddash A_{k}\circleddash\{a_{k}\}$ to locate its position.
2.2 Modeling of the Workspace
@@
We further establish a coordinate system $\mathbb{O}$ by picking an arbitrary point $O\in L$ as the origin, and denote the lattice points adjacent to the origin as $p_{1},p_{2},...,p_{6}$, labeled clockwise. We denote the line going through $p_{1},p_{2}$ as $l_{1}$ and the line going through $p_{4},p_{5}$ as $l_{2}$. Notice that $l_{1}$ divides $\mathbb{R}^{2}$ into two open half-plane [^2] An open half-plane does not include the line that induces the half-plane.. One of the open half-planes does not contain $l_{2}$, and we denote it as $H_{1}$. Similarly, the open half-plane induced by $l_{2}$ that does not contain $l_{1}$ is denoted as $H_{2}$. The idle robots are to be placed in $H_{2}$ and we call $H_{2}$ the idle half-plane. The robots will form the desired shape in $H_{1}$ and we call $H_{1}$ the formation half-plane. The buffering region located between $l_{1}$ and $l_{2}$ ensures no physical interaction between idle robots and the robots in $S$.
@@ 2.3 The Concept of Epoch
We denote the $n$th moving robot as $u_{n}$, which is referred as the active robot of the epoch $T=n$, and the rest of the robots as the inactive robots of the epoch $T=n$.
@@
The active robot $u_{n}$ is the only moving robot of the sub-epoch $T=n^{-}$ and the remaining robots are non-moving robots. There is no moving robot in the sub-epoch $T=n^{+}$ and all the robots are non-moving robots. For the case when $T=0$, there is only one sub-epoch $T=0^{+}$ where all the robots are in their initial configuration and no robot is moving.
@@
2.4 Modeling of the Robots
A lattice point $x\in L$ is occupied by a robot $u$ at time $t$ if $p_{t}(u)=x$; otherwise, it is unoccupied. If a vertex of ribbon $R$ is occupied by a robot $u$, we say the robot $u$ is located on the ribbon $R$. Given a ribbon $R$, we say the ribbon is filled if all the vertices of $R$ are occupied; the ribbon is empty if no vertex of $R$ is occupied; a ribbon is half-filled if it is neither filled nor empty. We say a robot belongs to a ribbon if the robot occupies one of the vertex of the ribbon. The robot occupying the leader vertex of a ribbon is the leader robot of the ribbon.
@@ 3.1 Ribbon Structure
    by considering $(p_{2},p_{1})$ as a ribbon, all ribbons form a tree structure rooted at $(p_{2},p_{1})$. For any two ribbons $R_{i}$ $R_{j}$ of the tree, we use $R_{i}\rightarrow R_{j}$ to denote that there exists a chain of ribbons $(R_{i},R_{i+1},\ldots,R_{j-1},R_{j})$ where $R_{i}$ is a parent ribbon of $R_{i-1}$,…,$R_{j-1}$ is a parent ribbon of $R_{j}$.
Then we check whether the ribbon structure satisfies the following conditions:
@@ 3.2 Planning of Movement Sequences
    The ribbon order is a total order of the lattice points of a ribbon $R$ such that for all different $x,y$ of ribbon $R$, $x<y$ if there is a directed path from $y$ to $x$.
@@
    The tree order is a partial order of the ribbons of $S$ such that for all different ribbons $R_{i},R_{j}$, $R_{i}<R_{j}$ if there exists a path $l$ in the tree structure induced by ribbons, while $l$ starts from the root $R_{0}$, passes $R_{i}$ and ends at $R_{j}$. A complete tree order is a total order that extends from the tree order [^7] A partial order extends to a total order by the Szpilrajn extension theorem..
@@
    The idle ribbon order is a total order of the lattice points of an idle ribbon $IR$ such that for all different $x,y$ of idle ribbon $IR$, $x<y$ if there is a directed path from $y$ to $x$.
@@
    The idle tree order is a total [^8] Notice that the tree of idle ribbons has a single branch. order of the idle ribbons such that for all different idle ribbons $IR_{i},IR_{j}$, $IR_{i}<IR_{j}$ if there exists a path $l$ in the tree structure induced by idle ribbons, while $l$ starts from $IR_{0}$, passes $IR_{i}$ and ends at $IR_{j}$.
@@
The re-activation sequence $s_{r}[k]$ for $1\leq k\leq N$ is obtained recursively such that each element $s_{r}[k]$ is the largest lattice point in $S$ (w.r.t the ribbon order) on the largest ribbon (w.r.t the complete tree order) which has not been included in ${s_{r}}_{[1:k-1]}$. The re-assembly sequence $t_{r}[k]$ for $1\leq k\leq N$ is generated recursively such that each element $t_{r}[k]$ satisfies:
@@
    if $s_{r}[k]$ is not among the first $C(R(s_{r}[k]))$ lattice points of the ribbon $R(s_{r}[k])$, $t_{r}[k]$ is the largest lattice point in $S\setminus D$ (w.r.t the ribbon order) on the largest ribbon (w.r.t the complete tree order) which has not been included in ${t_{r}}_{[1:k-1]}$.
@@ 3.3 Execute of the Algorithm
    Gradient-Formation: The root robot maintains a constant gradient value of 0. All the other robots choose the smallest gradient value among their neighboring robots, and set their own gradient value to be the minimum gradient value plus $1$.
@@
    Ribbon-Following: the moving robot orbits around a ribbon in the clockwise direction by maintaining distance $d$ to the closest robot located on the same ribbon. Please be reminded that $d=\frac{2r}{\sqrt{3}-1}$ is slightly larger than 2r. Thus, the collision can still be avoided while a robot performs ribbon-following movement within the shape $S$.
@@
Now we explain the execution of the algorithm step-by-step. Each robot $v$ continuously broadcast its current coordinates $p_{t}(v)$, its gradient value $g(v)$ and its epoch $T(v)$ to the robots in its neighborhood. Each robot $v$ will also receive the coordinates $p_{t}(u)$, the distance $d(u,v)$, the gradient value $g(u)$ and the local epoch $T(u)$ from all its neighboring robot $u\in N_{t}(v)$. Here we provide a flow diagram for the execution of the algorithm and the corresponding pseudocodes.
@@
    if $N+1\leq T\leq 2N$, the robot located at $s_{r}[T-N]$ starts ribbon-following until it is outside the shape $S$. Then the robot executes the edge-following algorithm until it stops at $t_{r}[T-N]$. After each robot stops, the epoch count is increased by 1 until $T=2N$, and the algorithm stops. In the subtractive stage, extra robots exit the hole $D$ and assemble in the idle half-plane.

### v3

@@ Abstract
Abstract: The pattern formation task is commonly seen in a multi-robot system. In this paper, we study the problem of forming complex shapes with functionally limited mobile robots, which have to rely on other robots to precisely locate themselves. The goal is to decide whether a given shape can be filled by a given set of robots; in case the answer is yes, to complete a shape formation process as fast as possible with a minimum amount of communication. Traditional approaches either require global coordinates for each robot or are prone to failure when attempting to form complex shapes beyond the capability of given approaches - the latter calls for a decision procedure that can tell whether a target shape can be formed before the actual shape-forming process starts. In this paper, we develop a method that does not require global coordinate information during the execution process and can effectively decide whether it is feasible to form the desired shape. The latter is achieved via a planning procedure that is capable of handling a variety of complex shapes, in particular, those with holes, and assigning a simple piece of scheduling information to each robot, facilitating subsequent distributed execution, which does not rely on the coordinates of all robots but only those of neighboring ones. The effectiveness of our shape-forming approach is vividly illustrated in several simulation case studies.
@@ 1 Introduction
Multi-Robot Systems (MRS), as suggested by the name, involves the study of groups of heterogeneous or homogeneous autonomous robots tackling a wide range of tasks with varying characteristics in the same environment [1]. Among the various tasks assigned to a group of autonomous robots, the deployment task has received significant attention since it is usually used to set up the initial configuration for other downstream tasks such as exploration, surveillance and cooperative manipulation. The pattern formation task, as a variant of the robot deployment tasks, aims to assemble a collection of robots into proper positions that resemble a specific pattern when viewed globally [2]. Multiple research studies have been performed on the pattern formation using homogeneous autonomous robots.
In [3], Stoy and Kasper proposed a shape formation algorithm utilizing a porous scaffold and established two types of gradients: the wander gradient, which is generated from wandering robots outside the desired shape, and the hole gradient, which is generated from the unfilled spaces inside the desired shape. A wandering robot descends the hole gradient and ascends the wander gradient, resulting in the robots inside the shape moving away from the boundary and propagating the unfilled spaces toward the boundary. In such a way, this algorithm is able to reconfigure a random shape into a desired non-hollow shape.
@@
In [10], Derakhshandeh et al. designed a shape formation algorithm based on the triangular expansion behavior of programmable matters. The algorithm can be adapted to any complex shape composed of equilateral triangles. However, similar to Yang’s approach, without global information on coordinates, the robots may not be able to form the holes correctly.
Current methods either lack the ability to form complex shapes involving holes or require global coordinates for accurate hole forming. In this article, we propose a pattern formation method, the add-subtract algorithm, to handle formation shapes with holes. In our algorithm, information on global coordinates is not required by each robot, as they could locate themselves based on the position of neighboring robots. The algorithm is inspired by the ribbon idea introduced in the additive self-assembly algorithm [4]. In our algorithm, we assume a grid-based workspace and choose the hexagonal lattice due to its compactness and suitability for disk-shaped robots. Several studies have explored hexagonal lattice configurations. K. Tomita et al. developed a nucleation method using hexagon-shaped robots [11]; M. De Rosa et al. defined hole motion in a hexagonal lattice for modular robot shape sculpting [12]; and M. Rubenstein et al. utilized hexagonal arrangements to organize idle robots [4]. Building on this foundation, we propose a systematic method to partition the lattice points inside the desired shape into layers (referred to as ribbons) and plan the movement sequences for each individual robot. During the execution phase, robots first occupy all the lattice points, including the holes, layer by layer, and then the extra robots move out of the holes in reverse order (starting from the outermost layer). Compared with other existing methods, our method bears the following distinct features:
@@
    the planning procedure can assist users to decide whether a complex shape, in particular, those with holes, can be formed by robots with limited localization capabilities; and
@@
    in the case the given shape is feasible, a concrete moving sequence will be generated, and each robot will be assigned with some simple piece of information that does not rely on the coordinates of all robots but only those of neighboring ones, thus facilitating distributed execution; and
@@
    the distributed execution phase can be further enhanced with fault tolerance capabilities via fast re-planning, thanks to our effective planning procedure, making the method robust to uncertainties in implementation.
The remainder of the paper is organized as follows. First, the problem is introduced and formulated in section 2. The add-subtract algorithm is introduced and analyzed in section 3. To ensure a smooth presentation flow, while we present the outline of the proof of correctness in section 4, details of all proofs are presented in the Appendices. Simulation results are given in section 5. Conclusions are drawn in section 6 together with a discussion of future research directions. For the pseudocode of the algorithm, please refer to appendix A. For the proof of the properties of the ribbon structure and the proof of correctness of the algorithm can be found in appendix B and appendix C.
@@ 2 Problem Formulation
In this section, we give a formal definition of the problem (section 2.1) and introduce the setup of the system (section 2.2). Finally, we introduce the concept of epoch used in the algorithm (section 2.3).
@@ 2.1 Problem Description
In this paper, we study the problem of filling up a given shape with a given collection of small mobile robots. The shape may contain holes, and each mobile robot has a limited localization capability, forcing them to rely on neighboring robots to precisely determine its current location. Due to this limitation, during the shape-filling process, robots must follow a certain releasing sequence to ensure that each newly released robot can rely on previously released and positioned robots to precisely locate itself, before moving into a designated position in the shape. We assume that each robot has sufficient memory to store several pieces of critical information, such as the releasing sequence, which encoded the boundaries of the shape and internal holes. In addition, we assume that each robot can only communicate with robots within a neighborhood whose maximum radius is known in advance. Next, we provide more details of the system setup.
@@
We consider a user-specified shape $S_{0}$ defined in the Euclidean coordinate system in $\mathbb{R}^{2}$ which is topologically homeomorphic to an open disk in $\mathbb{R}^{2}$. We consider a set $D_{0}$ of user-specified Holes defined in the Euclidean coordinate system in $\mathbb{R}^{2}$, each of which is topologically homeomorphic to a collection of closed disks in $\mathbb{R}^{2}$ and contained inside $S_{0}$. With a slight abuse of notation, we use $D_{0}$ to denote the union of all such holes. Thus, $D_{0}$ is a subset of $S_{0}$.
@@
    all the lattice points in $S$ are connected (in the sense that there is a path between any pair of lattice points in $S$ through adjacent lattice points in $S$); and
@@
    For each prefix subsequence $(a_{1},b_{1},1),\ldots,(a_{k},b_{k},k)$, $1\leq k\leq M$, let $A_{k}=\{a_{1},\ldots,a_{k-1}\}$, $B_{k}=\{b_{1},\ldots,b_{k-1}\}$, and $\overline{B_{k}}\subseteq B_{k}$ be the set of stopping positions that have been reached by the corresponding moving robots. Then the starting position of the $k$th robot $a_{k}\in I\oplus\overline{B_{k}}\circleddash A_{k}$ and the robot can only relay on robots centered at $I\oplus\overline{B_{k}}\circleddash A_{k}\circleddash\{a_{k}\}$ to locate its position before stopping at $b_{k}$.
2.2 Setup of the System
@@
We further establish a coordinate system $\mathbb{O}$ by picking an arbitrary point $O\in L$ as the origin and denote the lattice points adjacent to the origin as $p_{1},p_{2},...,p_{6}$, labeled clockwise. We denote the line going through $p_{1},p_{2}$ as $l_{1}$ and the line going through $p_{4},p_{5}$ as $l_{2}$. Notice that $l_{1}$ divides $\mathbb{R}^{2}$ into two open half-planes [^2] An open half-plane does not include the line that induces the half-plane.. One of the open half-planes does not contain $l_{2}$, and we denote it as $H_{1}$. Similarly, the open half-plane induced by $l_{2}$ that does not contain $l_{1}$ is denoted as $H_{2}$. The idle robots are to be placed in $H_{2}$ and we call $H_{2}$ the idle half-plane. The robots will form the desired shape in $H_{1}$ and we call $H_{1}$ the formation half-plane. The buffering region located between $l_{1}$ and $l_{2}$ ensures no physical interaction between idle robots and the robots in $S$.
A lattice point $x\in L$ is occupied by a robot $u$ at time $t$ if $p_{t}(u)=x$; otherwise, it is unoccupied. If a vertex of ribbon $R$ is occupied by a robot $u$, we say the robot $u$ is located on the ribbon $R$. Given a ribbon $R$, we say the ribbon is filled if all the vertices of $R$ are occupied; the ribbon is empty if no vertex of $R$ is occupied; a ribbon is half-filled if it is neither filled nor empty. We say a robot belongs to a ribbon if the robot occupies one of the vertices of the ribbon. The robot occupying the leader vertex of a ribbon is the leader robot of the ribbon.
@@ 2.3 The Concept of Epoch
We denote the $n$th moving robot as $u_{n}$, which is referred to as the active robot of the epoch $T=n$, and the rest of the robots as the inactive robots of the epoch $T=n$.
@@
The active robot $u_{n}$ is the only moving robot of the sub-epoch $T=n^{-}$, and the remaining robots are non-moving robots. There is no moving robot in the sub-epoch $T=n^{+}$, and all the robots are non-moving robots. For the case when $T=0$, there is only one sub-epoch $T=0^{+}$ where all the robots are in their initial configuration and no robot is moving.
@@ 3.1 Ribbon Structure
    by considering $(p_{2},p_{1})$ as a ribbon, all ribbons form a tree structure rooted at $(p_{2},p_{1})$. For any two ribbons $R_{i}$ $R_{j}$ of the tree, we use $R_{i}\twoheadrightarrow R_{j}$ to denote that there exists a chain of ribbons $(R_{i},R_{i+1},\ldots,R_{j-1},R_{j})$ where $R_{i}$ is a parent ribbon of $R_{i-1}$,…,$R_{j-1}$ is a parent ribbon of $R_{j}$.
Then, we check whether the ribbon structure satisfies the following conditions:
@@ 3.2 Planning of Movement Sequences
    The ribbon order is a total order of the lattice points of a ribbon $R$ such that for all different $x,y$ of ribbon $R$, $x<y$ if $y\rightarrow x$.
@@
    The tree order is a partial order of the ribbons of $S$ such that for all different ribbons $R_{i},R_{j}$, $R_{i}<R_{j}$ if $R_{i}\twoheadrightarrow R_{j}$. A complete tree order is a total order that extends from the tree order [^7] A partial order extends to a total order by the Szpilrajn extension theorem..
@@
    The idle ribbon order is a total order of the lattice points of an idle ribbon $IR$ such that for all different $x,y$ of idle ribbon $IR$, $x<y$ if $y\rightarrow x$.
@@
    The idle tree order is a total [^8] Notice that the tree of idle ribbons has a single branch. order of the idle ribbons such that for all different idle ribbons $IR_{i},IR_{j}$, $IR_{i}<IR_{j}$ if $IR_{i}\twoheadrightarrow IR_{j}$
@@
The re-activation sequence $s_{r}[k]$ for $1\leq k\leq N$ is obtained recursively such that each element $s_{r}[k]$ is the smallest lattice point in $S$ (w.r.t the ribbon order) on the largest ribbon (w.r.t the complete tree order) which has not been included in ${s_{r}}_{[1:k-1]}$. The re-assembly sequence $t_{r}[k]$ for $1\leq k\leq N$ is generated recursively such that each element $t_{r}[k]$ satisfies:
@@
    if $s_{r}[k]$ is not among the first $c(R(s_{r}[k]))$ lattice points of the ribbon $R(s_{r}[k])$, $t_{r}[k]$ is the smallest lattice point in $S\setminus D$ (w.r.t the ribbon order) on the largest ribbon (w.r.t the complete tree order) which has not been included in ${t_{r}}_{[1:k-1]}$.
@@ 3.3 Execute of the Algorithm
    Gradient-Formation: The root robot maintains a constant gradient value of 0. All the other robots choose the smallest gradient value among their neighboring robots and set their own gradient value to be the minimum gradient value plus $1$.
@@
    Ribbon-Following: the moving robot orbits around a ribbon in the clockwise direction by maintaining distance $d$ to the closest robot located on the same ribbon. Please be reminded that $d=\frac{2r}{\sqrt{3}-1}$ is slightly larger than 2r. Thus, the collision can still be avoided while a robot performs the ribbon-following movement within the shape $S$.
@@
Now, we explain the execution of the algorithm step-by-step. Each robot $v$ continuously broadcasts its current coordinates $p_{t}(v)$, its gradient value $g(v)$ and its epoch $T(v)$ to the robots in its neighborhood. Each robot $v$ will also receive the coordinates $p_{t}(u)$, the distance $d(u,v)$, the gradient value $g(u)$ and the local epoch $T(u)$ from all its neighboring robot $u\in N_{t}(v)$. Here, we provide a flow diagram for the execution of the algorithm and the corresponding pseudocodes.
@@
    if $N+1\leq T\leq 2N$, the robot located at $s_{r}[T-N]$ starts ribbon-following until it is outside the shape $S$. Then, the robot executes the edge-following algorithm until it stops at $t_{r}[T-N]$. After each robot stops, the epoch count is increased by 1 until $T=2N$, and the algorithm stops. In the subtractive stage, extra robots exit the hole $D$ and assemble in the idle half-plane.